# Matching Non-Identical Objects

### v1

@@ Abstract
Not identical but similar objects are everywhere in the world. Examples include four-legged animals such as dogs and cats, cars of different models, akin flowers in various colors, and countless others. In this study, we address a novel task of matching such non-identical objects. We propose a simple weighting scheme of descriptors that enhance various sparse image matching methods, which are originally designed for matching identical objects captured from different perspectives, and achieve semantically robust matching. The experiments show successful matching between non-identical objects in various cases including domain shift. Further, we present a first evaluation of the robustness of the image matching methods under common corruptions, which is a sort of domain shift, and the proposed method improves the matching in this case as well.
@@ 1 Introduction
Figure 1: Non-identical object matching is achieved densely and consistently by adding our method between SuperPoint [7] and LightGlue [17]. Object correspondences can be found even when the design is different as in (a), the classes are discrepant as in (b), and there are several similar objects as in (c). Objects are robustly matched to each other even with domain shifts in pictures and photos as in (d) and corruption as in (e) and (f).
We humans are able to find fine-grained visual correspondences between two objects even when they look similar but not identical or when one is sketched and another appears in a photo. Thanks to this ability, for example, we can identify a cat in gloomy backstreet from a photo taken in a bright room or assemble components to build a shelf based on its perspective drawing. In computer vision words, this needs fine-grained matching between feature points of two objects under image corruptions, class shift, and domain shift. The classical image matching task [30, 36, 17, 11] only addresses a special case; objects to be matched are assumed to be identical and captured from different perspectives. Recent Transformer-based dense matching methods show a great performance on such a matching task [36, 26, 48, 14, 8, 46, 4, 45], and there are also light-weight sparse matching methods, which run in real time with a slight cost in accuracy [30, 3, 44, 27, 23, 11, 19]. However, they do not perform well once the identicality assumption is not violated and/or objects are captured in harsh environments or different domains (cf. Fig. 2).
@@
In this study, we address a novel task of non-identical object matching, where objects to be matched are not necessarily identical. For example, the objects to be matched can be clocks of different designs or animals of different species. Further, the non-identicality naturally induces the necessity of robustness against image corruptions (e.g., blurs and brightness change) and domain shifts (e.g., photo v.s. drawing) because two objects may be captured in different environments and domains, as shown in Fig. 1. A crucial challenge of achieving non-identical object matching is that we cannot simply resort to a supervised learning framework because of the ambiguity of non-identicality; there is no ground-truth matching between non-identical objects. At least, to our knowledge, there are no such datasets. Interestingly, we discovered a simple plug-and-play weighting module can extend off-the-shelf sparse matching methods to the non-identical object matching. Namely, given a pair of images, sparse matching methods first perform keypoint detection and then feed the extracted descriptors to a feature matching model to determine the matching of the keypoints. Our weighting module can be inserted before the feature matching model to weights the descriptors using semantic information. Particularly, we introduce an object detector to identify objects and compute a Grad-CAM heatmap [32] to obtain weights. Previously, because of the task definition, matching methods only exploit low-level features, such as colors, edges, and texture, which are all susceptible to ambiguities and perturbations appearing in non-identical object matching task. The low-level features are still essential to realize fine-grained matching, but the proposed module makes it robust using high-level features.
@@
Our experiments demonstrate that this simple approach significantly boosts the robustness of sparse matching methods, including X and Y, in two tasks: non-identical object matching and robust image matching. In the former task, we evaluate the matching results with the image pairs (a) of the same class, (b) of different classes, (c) with domain shift, and (d) of the same appearance. The sparse matchers alone resulted in scattered mismatching between non-corresponding areas, but with the addition of our method, fine-grained and consistent matching is achieved. In the latter task, we evaluate the robustness of matching on the MegaDepth-1500 [36] dataset under common corruptions [13] in terms of the AUC score of standard relative pose estimation. The proposed method increases the AUC scores with various types of common corruptions, and exceeds a dense matching method LoFTR [36] in the average case.
@@
  • We propose non-identical object matching, a novel task of matching similar but unnecessarily identical objects in images even under image corruptions and/or between different domains. This task can be considered a generation of the classical image matching task.
@@
  • We present a first evaluation of the robustness in image matching under common corruptions. By adding corruptions to both input images, the robustness of corruptions itself is measured. By adding them to only one of the input images, the robustness against a domain shift is measured. For both cases, the proposed method is shown useful.
@@ 2 Related Work
Image matching [30, 36, 17, 11] is a fundamental task in computer vision with various applications [31, 22, 5, 25, 21, 42, 37, 35] that involves finding the fine-grained correspondence between two images. Many image matching methods consist of two stages. First, a keypoint detector extracts collections of keypoints, which are pairs of positions and the local feature descriptors. Then, feature matchers find a match between them.
@@
Classically, keypoint detection was down using hand-crafted local features such as SIFT [18] and ORB [29]. Recently, learning-based methods have been developed [7, 12, 10, 40]. Particularly, SuperPoint [7] and DISK [38] are two popular examples. The former is based on convolutional neural networks, and the latter is based on reinforcement learning. There is also a fully differentiable and lightweight method with a minimal model configuration [10]. Unsupervised keypoint detection based on a text-to-image diffusion model [12] also exists. In [40], a flexible technique, which can be combined with other keypoint detectors, is proposed to extract keypoints for robustly performing a matching under illumination changes and repeating texture patterns.
@@
Feature matchers [43] find a fine-grained correspondence between images. The feature matchers are categorized into two types based on the density of matching points: sparse matchers and dense matchers. The former runs faster, while the latter gives dense and high accuracy. The focus of our study lies in sparse matchers. A seminal sparse matcher, SuperGlue [30], addresses the partial assignment problem for matching by integrating the Transformer attention mechanism [39] with optimal transport [24]. The state-of-the-art sparse matcher, LightGlue [17], improves SuperGlue by introducing a hierarchical matching structure, achieving high matching accuracy and real-time processing simultaneously. GlueStick [23] utilizes line segments for structural features. In contrast to sparse matchers, dense matchers [36, 26, 48, 14, 8, 46, 4, 45] peruse matching accuracy rather than speed. Dense matchers input an image and infer matching using all pixels as keypoints. Therefore, it does not require a keypoint detector so it is also called a detector-free model. LoFTR [36] proposed the first transformer-based method for dense matching, which achieves highly dense matching. Recent studies such as ASpanFormer [4], which controls the attention span, and ASTR [45], which focuses on the role of each pixel.
The aforementioned methods are all designed for image matching and do not perform well in our new task, where objects to be matched do not necessarily appear the same and/or images are corrupted (cf. Fig. 6). There are a few studies that aim at robust matching in particular cases. For example, DarkFeat [11] address matching in low-light conditions, and SAM [19] peruses robustness against large viewpoint change and illumination change. As for such image corruptions, we consider fifteen types of common corruptions [13], including large brightness change and contrast change.
@@ 3 Non-Identical Object Matching
This study introduces non-identical object matching, matching between unnecessarily identical objects in two images. For example, a dog and a cat are not identical, but still have similar structures such as two eyes, four legs, and one tail. The same mass-produced products, such as cars of the same model, have the same appearance but still not identical (e.g., owned by different people). Thus, they can appear in images with largely different backgrounds and lighting conditions.
@@ 3.1 Task formulation
In the followings, we present a more formal problem setup based on a standard matching pipeline. Note that it is essentially hard to completely formalize the non-identical object matching because of the ambiguity of non-identicality, so our formalization is slightly conceptual.
@@
Let ${\mathcal{X}}$ be the image domain. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, let $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i=1}^{n_{A}}$ be $n_{A}$ pairs of keypoint $\bm{p}_{i}^{A}\in{\mathbb{R}}^{2}$ and its descriptor $\bm{d}_{i}^{A}\in[0,1]^{d}$ of image $\bm{x}_{A}$. We define $D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i=1}^{n_{B}}$ for image $\bm{x}_{B}$ similarly. Let ${\mathcal{L}}(\,\cdot\,)$ be a (conceptual) matching loss function. The non-identical object matching can be formalized as follows.
$\min\quad$ (${\mathcal{L}}(\{m_{kl}\}_{k=1,\ldots,n_{A},l=1,\ldots,n_{B}})$ (1)
@@
The matching coefficient $m_{kl}=1$ indicates that the $k$-th keypoint in image $\bm{x}_{A}$ and the $l$-th keypoint in image $\bm{x}_{B}$ is matched and otherwise $m_{kl}=0$. Each keypoint in one image has at most one matched point in the other image. This condition is implemented by the last two constraints. In the image matching task, where an identical object is captured from different angles in two images, the matching loss (${\mathcal{L}}(\{m_{kl}\}_{k,l})$ measures the discrepancy between a camera pose computed the matching $\{m_{kl}\}_{k,l}$ and the ground truth.
Our task, non-identical object matching, covers a broader class of matching. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$ that contain (potentially non-identical) objects to be matched, image $\bm{x}_{i}$ can be characterized by several attributes $(o_{i},y_{i},P_{i},D_{i})$ for $i\in\{1,2\}$. If $o_{A}=o_{B}$, then the objects to be matched are identical. If $o_{A}\equiv o_{B}$, they are the same in appearance but not necessarily identical (e.g., two cars of the same model). If $y_{A}=y_{B}$, they belong to the same class (e.g., the car class). If $P_{A}=P_{B}$, then they are captured from the same camera pose. If $D_{A}=D_{B}$, they are captured in the same domain (e.g., the drawing domain).
The classical image matching corresponds to the case with $o_{A}=o_{B}$ (consequently, $o_{A}\equiv o_{B}$ and $y_{A}=y_{B}$), $P_{A}\neq P_{B}$, and $D_{A}=D_{B}$. In words, it considers identical objects captured from different camera poses in the same domain. In contrast, non-identical object matching encompasses $y_{A}\neq y_{B}$ (consequently, $o_{A}\neq o_{B}$ and $o_{A}\not\equiv o_{B}$), $P_{A}\neq P_{B}$, and $D_{A}\neq D_{B}$. Note that we practically assume $o_{A}$ and $o_{B}$ to be similar in some sense. Below, we highlight two important special cases. For both cases, one may assume $P_{A}\neq P_{B}$.
Figure 2: Failure cases of non-identical object matching by SuperPoint [7] as a keypoint detector and LightGlue [17] as the state-of-the-art matcher. (a) The images are matched with backgrounds that have no correspondence. (b) The texture of the bear and the background are matched. (c) The shape of the corruption called frost is matched, because the matcher does not know the existence of the object. (d) There is a large difference in the matches between the right and left eyes. This is because the matcher does not consider the semantics of the object.
@@
Four-legged animals, such as tigers and cats, are objects of different classes, but have similar structures. They both have two eyes, four legs, and one tail. We humans can find a part-by-part matching between them, even when there are differences in color, texture, and shape. However, it is not trivial how one can design a matching methods to find a fine-graind correspondence while being robust against the difference in details. If such a matching is realized, for example, one may track the running motions of unknown four-legged animals based on those of known ones.
Figure 3: Pipeline of the matching. The keypoint detention and feature matching are done by off-the-self models. The proposed weighting module computes the heatmap scores of objects and weights the descriptors with these semantic information.
@@
We humans can know the flower in the painting on the well depicts that in the garden. This is a robust fine-graind matching of an identical object in different domains. Typical domain shifts include drawings v.s. photos and clean images v.s. corrupted images. For example, suppose that $\bm{x}_{A}$ is a photo and $\bm{x}_{B}$ is a painting. Compared to photos, drawings have less colors, textures, and shadows. Thus, while the descriptors of the photo $\bm{x}_{A}$ may reflect the color or texture, those of the drawing $\bm{x}_{B}$ do not, which can lead to an absent of correspondence if the feature matcher is not sufficiently robust. If robust matching between different domains is realized, for example, one can design a system that guides a worker to assemble parts from the first-person video by comparing the parts with those in a blueprint.

### v3

@@ Abstract
Not identical but similar objects are ubiquitous in our world, ranging from four-legged animals such as dogs and cats to cars of different models and flowers of various colors. This study addresses a novel task of matching such non-identical objects at the pixel level. We propose a weighting scheme of descriptors that incorporates semantic information from object detectors into existing sparse image matching methods, extending their targets from identical objects captured from different perspectives to semantically similar objects. The experiments show successful matching between non-identical objects in various cases, including in-class design variations, class discrepancy, and domain shifts (e.g., photo–drawing and image corruptions).
@@ 1 Introduction
We humans are able to find fine-grained and robust visual correspondences between two objects even when they are not necessarily identical or when they are captured in different situations. Thanks to this ability, for example, we can find out our missing cats in a gloomy backstreet by remembering them basking in the sun or assembling components to build a shelf based on their perspective drawing. In computer vision words, this needs fine-grained matching of feature points between two objects even with class discrepancy and domain shift (e.g., image corruptions).
The abovementioned matching has been partially addressed in the literature. In the image matching [49, 56, 27, 17], objects to be matched are assumed identical and captured from different perspectives. This assumption also enables supervised training using dense correspondence obtained from 3D reconstruction from image pairs. Recently, Transformer-based dense matching methods have shown great performance [56, 45, 74, 23, 11, 72, 5, 67], and there are also light-weight sparse matching methods, which run in real time with a slight cost in accuracy [49, 66, 46, 38, 17, 30, 22]. In contrast, semantic correspondence models [70, 32, 7] focus on matching semantically similar parts between objects (e.g., tires of two bikes). They are trained on datasets with sparse part-to-part correspondence annotation and thus do not offer fine-grained matching.
Figure 1: Non-identical object matching by our plug-and-play module incorporated with the sparse image matching pipeline (here, SuperPoint [10] for keypoint detection and LightGlue [27] for matching). (a) design differences, (b) class discrepancy, (c) similar objects, and domain shifts such as (d) drawing v.s. photo and (e, f) different image corruptions.
In this study, we generalize image matching and semantic correspondence to non-identical object matching, where objects to be matched are not necessarily identical (see Fig. 1). For example, the object pair can be street clocks of different designs or four-legged animals of different species. Further, the non-identicality naturally requires robustness, e.g., against image corruptions (e.g., blurs and brightness change) and domain shifts (e.g., photo v.s. drawing) because two objects may be captured in different environments and domains. A crucial challenge of achieving non-identical object matching is that we cannot straightforwardly resort to the supervised learning framework because of the ambiguity of non-identicality; there is no ground-truth matching between non-identical objects.
@@
We propose a simple plug-and-play weighting module that extends off-the-shelf sparse matching methods to non-identical object matching. Given a pair of images, existing sparse matching methods first perform keypoint detection and then feed the extracted descriptors to a feature matching model to determine the matching of the keypoints. Our weighting module is inserted before the feature matching model and weights the descriptors using semantic information. Particularly, we introduce an object detector to identify objects and compute a Grad-CAM heatmap [51] to obtain weights. Previously, because of the task definition, matching methods only exploit low-level features, such as colors, edges, and texture, which are all susceptible to ambiguities and perturbations appearing in the non-identical object matching task (see Fig. 2). The low-level features are still essential to realize fine-grained matching, but the proposed module makes it robust using high-level features. Further, to avoid emphasizing irrelevant objects when an image contains more than one object, we propose Class Similarity Adapter, which determines the most relevant object pairs. We leverage the class labels from an object detector and determine object pairs based on their semantic similarity in the text embedding space.
@@
Our experiments demonstrate that our approach significantly boosts the robustness of sparse matching methods, including LightGlue [27] and GlueStick [38], in two tasks: non-identical object matching and robust image matching. In the former task, we evaluate the matching results with various image pairs (a) of the same class, (b) with class discrepancy, (c) with domain shift, and (d) of the same appearance. The sparse matchers alone result in scattered mismatching between non-corresponding areas, but combined with our method, fine-grained and consistent matching is achieved. In the latter task, we evaluate the robustness of matching on the MegaDepth-1500 [56] dataset under common corruptions [20] in terms of the AUC score of standard relative pose estimation. Our method outperforms the state-of-the-art sparse matcher under various types of image corruptions. The average AUC of our method even exceeds that of dense matchers, LoFTR [56] and Efficient LoFTR [64], which pursue matching quality at the cost of speed.
@@
  • We tackle non-identical object matching, a novel task of matching similar but unnecessarily identical objects in images, even under image corruptions and/or between different domains, which can be considered a generation of classical image matching and semantic correspodense.
@@
  • We present a first evaluation of the robustness in image matching under common corruptions. By adding corruptions to both input images, the robustness against corruptions itself is measured. By adding them to only one of the input images, the robustness against a domain shift is measured. For both cases, our method is shown useful.
@@ 2 Related Work
Figure 2: Failure cases of non-identical object matching using the state-of-the-art matcher, LightGlue [27], and the semantic correspondence model, GeoAware-SC [70]. In semantic correspondence models, some points indicating object parts must be designated during inference to find correspondences, and automating this point selection is essential for feature point matching. Comparison of feature matching by GeoAware-SC using two methods: one that randomly extracts 50 points and the other that adaptively extracts 50 points based on the GradCAM [51] heatmap of YOLOv7 [62]. Red boxes indicate mismatches due to background or object-like textures. Orange boxes indicate semantically corresponding regions, where most points align at a part level but not at a pixel level. For example, in (d), mismatches occur between the tip of the nose of the cheetah and the nostrils of the husky.
Image matching [49, 56, 27, 17] is a fundamental task in computer vision with various applications [50, 35, 8, 43, 34, 63, 57, 55] that involves finding the fine-grained correspondence between two images. Many image matching methods have two stages. A keypoint detector first extracts keypoints with their positions and local feature descriptors, and then feature matchers find correspondences between them.
@@
Classically, keypoint detection was done using hand-crafted local features such as SIFT [28] and ORB [48]. Recently, learning-based methods have been developed [10, 18, 14, 61, 41, 12]. Particularly, SuperPoint [10] and DISK [59] are two popular examples. The former is based on convolutional neural networks, and the latter is based on reinforcement learning. There is also a fully differentiable and lightweight method with a minimal model configuration [14]. Unsupervised keypoint detection based on a text-to-image diffusion model [18] also exists.
@@
Feature matchers [65, 68] find a fine-grained correspondence between images. The feature matchers are categorized into two types based on the density of matching points: sparse matchers and dense matchers. The former runs faster, while the latter gives dense and high accuracy. The focus of our study lies in sparse matchers. A seminal sparse matcher, SuperGlue [49], addresses the partial assignment problem for matching by integrating the Transformer attention mechanism [60] with optimal transport [40]. The state-of-the-art sparse matcher, LightGlue [27], improves SuperGlue by introducing a hierarchical matching structure, achieving high matching accuracy and real-time processing simultaneously. GlueStick [38] utilizes line segments for structural features. In contrast to sparse matchers, dense matchers [56, 45, 74, 23, 11, 72, 5, 67] peruse matching accuracy rather than speed. Dense matchers input an image and infer matching using all pixels as keypoints. Therefore, it does not require a keypoint detector, so it is also called a detector-free model. LoFTR [56] proposed the first transformer-based method for dense matching, which achieves highly dense matching. Recent studies [36, 71, 29] such as Efficient LoFTR [64], enabling fast and accurate matching by token aggregation, ASpanFormer [5], which controls the attention span, and ASTR [67], which focuses on the role of each pixel.
Semantic correspondence models.
Semantic Correspondence [15, 16, 33, 6, 2] is a task of identifying a semantically similar point in the target image from the given point in the source image. Specifically, it enables the correspondence of parts, such as the mouth, nose, or right eye of dogs between the source and target images. The training of a model is conducted using a specialized dataset, such as SPair-71k [33] and PF-PASCAL [15], which employ sparsely annotated part-level correspondences as supervisory data. Recently, GeoAware-SC [70] has been proposed as a model that leverages pre-trained features from two foundational models, Stable Diffusion [47] and DINOv2 [37], integrating both geometric and semantic information. This model significantly improves accuracy on major benchmarks for semantic correspondence, including the SPair-71k [33] and PF-PASCAL [15] datasets. In recent years, various models [39, 52, 69, 58, 21] have been studied. For example, CATs++ [7], which combines the local feature aggregation capacity of CNNs with the broad contextual comprehension of transformers; a model [32] employing feature maps from the reverse diffusion process in diffusion models; and unsupervised models [19] that learn semantic correspondence without requiring annotated data.
While feature matching and semantic correspondence are also the tasks of finding correspondences, our task (i.e., fine-grained matching between non-identical objects) is more challenging. First, it is hard to be addressed by the supervised learning framework because (i) manual annotation, as in semantic correspondence datasets is laborious for the pixel-level correspondence, and (ii) automatic annotation by 3D reconstruction, as in feature matching, which assumes an identical object viewed from different perspectives, cannot be done. Second, because non-identical objects generally appear with different lighting, image corruptions, or even image domains, the matching method should be designed to be robust to such diversity. As shown in Fig. 2, state-of-the-art feature matching methods and semantic correspondence models are susceptible to such variations. Several studies addressed robust matching with specific conditions, such as DarkFeat [17] for low-light and SAM [30] for viewpoint and illumination changes. We consider broader changes, including fifteen types of common corruptions [20] and several image domains (e.g., photos and sketches).
@@ 3 Non-Identical Object Matching
We introduce non-identical object matching, matching between unnecessarily identical objects in two images. For example, a dog and a cat are not identical but still have similar structures, such as two eyes, four legs, and one tail. The same mass-produced products, such as cars of the same model, have the same appearance but are still not identical (e.g., owned by different people) and can appear in images with largely different backgrounds and lighting conditions.
@@ 3.1 Task formulation
In the following, we present a more formal problem setup based on a standard matching pipeline. Note that it is essentially hard to completely formalize the non-identical object matching because of the ambiguity of non-identicality, so our formalization is slightly conceptual.
@@
Let $\mathcal{X}$} be the image domain. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$, let $D_{A}=\{(\bm{p}_{i}^{A},\bm{d}_{i}^{A})\}_{i=1}^{n_{A}}$ be $n_{A}$ keypoints in image $\bm{x}_{A}$, where $\bm{p}_{i}^{A}\in{\mathbb{R}}^{2}$ and $\bm{d}_{i}^{A}\in[0,1]^{d}$ are the position and the descriptor of $i$-th keypoint, respectively. We define $D_{B}=\{(\bm{p}_{i}^{B},\bm{d}_{i}^{B})\}_{i=1}^{n_{B}}$ for image $\bm{x}_{B}$ similarly. Let ${\mathcal{L}}(\,\cdot\,)$ be a (conceptual) matching loss function. The non-identical object matching can be formalized as follows.
$\min\quad$ ${\mathcal{L}}(\{m_{kl}\}_{k=1,\ldots,n_{A},l=1,\ldots,n_{B}})$ (1)
@@
The matching coefficient $m_{kl}=1$ indicates that the $k$-th keypoint in image $\bm{x}_{A}$ and the $l$-th keypoint in image $\bm{x}_{B}$ is matched and otherwise $m_{kl}=0$. Each keypoint in one image has at most one matched point in the other image. This condition is implemented by the last two constraints. In the image matching task, where an identical object is captured from different angles in two images, the matching loss ${\mathcal{L}}(\{m_{kl}\}_{k,l})$ measures the discrepancy between a camera pose computed the matching $\{m_{kl}\}_{k,l}$ and the ground truth.
Our task, non-identical object matching, covers a broader class of matching. Given two images $\bm{x}_{A},\bm{x}_{B}\in{\mathcal{X}}$ that contain (potentially non-identical) objects to be matched, image $\bm{x}_{I}$ can be characterized by several attributes $(o_{I},y_{I},P_{I},D_{I})$ for $I\in\{A,B\}$. If $o_{A}=o_{B}$, then the objects to be matched are identical. If $o_{A}\equiv o_{B}$, they are the same in appearance but not necessarily identical (e.g., two cars of the same model). If $y_{A}=y_{B}$, they belong to the same class (e.g., the car class). If $P_{A}=P_{B}$, then they are captured from the same camera pose. If $D_{A}=D_{B}$, they are captured in the same domain (e.g., the drawing domain).
The classical image matching corresponds to the case with $o_{A}=o_{B}$ (consequently, $o_{A}\equiv o_{B}$ and $y_{A}=y_{B}$), $P_{A}\neq P_{B}$, and $D_{A}=D_{B}$. In words, it considers identical objects captured from different camera poses in the same domain. In contrast, non-identical object matching encompasses $y_{A}\neq y_{B}$ (consequently, $o_{A}\neq o_{B}$ and $o_{A}\not\equiv o_{B}$), $P_{A}\neq P_{B}$, and $D_{A}\neq D_{B}$. Note that we practically assume $o_{A}$ and $o_{B}$ to be similar in some sense. Below, we highlight two important special cases, both assuming $P_{A}\neq P_{B}$.
Figure 3: Pipeline of the matching. The keypoint detection and feature matching are done by off-the-shelf models. The proposed weighting module computes the heatmap scores of objects and weights the descriptors with this semantic information.
@@
Four-legged animals, such as tigers and cats, are objects of different classes but have similar structures. They both have two eyes, four legs, and one tail. We humans can find a part-by-part matching between them, even with differences in color, texture, and shape. However, it is not trivial to design matching methods to find fine-grained correspondence while being robust to detail differences. If such a matching is realized, for example, one may track the running motions of unknown four-legged animals based on those of known ones.
@@
We humans can know the flower in the painting on the wall depicts that in the garden. This is a robust fine-grained matching of an identical object in different domains. Typical domain shifts include drawings v.s. photos and clean images v.s. corrupted images. For example, suppose that $\bm{x}_{A}$ is a photo and $\bm{x}_{B}$ is a painting. Compared to photos, drawings have fewer colors, textures, and shadows. Thus, while the descriptors of the photo $\bm{x}_{A}$ may reflect the color or texture, those of the drawing $\bm{x}_{B}$ do not, which can lead to an absent of correspondence if the feature matcher is not sufficiently robust. If robust matching between different domains is realized, for example, one can design a system that guides a worker to assemble parts from the first-person video by comparing the parts with those in a blueprint.